# Non-Equilibrium Steady States and Quantum Chaos in a three-site Driven-Dissipative Bose-Hubbard Chains base on Self-Consistent Mean-Field Approach

### v5

@@ VIII Quantum metric framework
For $s<0$, in which case the eigenvalues relative to the global decay $i\gamma/2$ are imaginary and thus corresponds to exponential growth and decay, $\zeta_{3}(t)=C_{1}\cosh(\sqrt{-s}t)+C_{2}\sinh(\sqrt{-s}t)+C_{3}\sim e^{\sqrt{-s}t}$, $A(t)\sim-|\vec{\lambda}|^{2}\zeta_{3}(t)\sim-e^{2\Gamma t}$. If the system enters the PT-broken phase where $\zeta_{3}(t)$ grows exponentially, $A(t)$ will evolve accordingly to ensure that the norm of the fluctuation vector $\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle=1$, which satisfies the time-dependent Schrodinger equation $i\partial_{t}|\mathbf{v}(t)\rangle=i\mathbf{L}_{\text{eff}}|\mathbf{v}(t)\rangle$. For $s<0$ we further have,
@@
$\langle u_{+}|\dot{\rho}(t)|u_{-}\rangle$ $=-\langle u_{+}|(\mathbf{L}_{\text{eff}}^{\dagger}\rho+\rho\mathbf{L}_{\text{eff}})|u_{-}\rangle=-[(\frac{\sqrt{-s}}{2}-\gamma/2)+(-\frac{\sqrt{-s}}{2}-\gamma/2)]\langle u_{+}|\rho(t)|u_{-}\rangle=\gamma\langle u_{+}|\rho(t)|u_{-}\rangle\implies\langle u_{+}|\rho(t)|u_{-}\rangle=\langle u_{+}|\rho(0)|u_{-}\rangle e^{\gamma t}$
@@ IX Appendix: Green’s Functions and Liouvillian Spectrum
The linear response of the system to a weak probe field is encoded in the local retarded Green’s function. In the frequency domain, this can be expressed using the resolvent of the effective Liouvillian superoperator

### v6

@@ VIII Quantum metric framework
For $s<0$, in which case the eigenvalues relative to the global decay $i\gamma/2$ are imaginary and thus corresponds to exponential growth and decay, $\zeta_{3}(t)=C_{1}\cosh(\sqrt{-s}t)+C_{2}\sinh(\sqrt{-s}t)+C_{3}\sim e^{\sqrt{-s}t}$, $A(t)\sim-|\vec{\lambda}|^{2}\zeta_{3}(t)\sim-e^{2\Gamma t}$. If the system enters the PT-broken phase where $\zeta_{3}(t)$ grows exponentially, $A(t)$ will evolve accordingly to ensure that the norm of the fluctuation vector $\langle\mathbf{v}(t)|\rho(t)|\mathbf{v}(t)\rangle=1$, which satisfies the time-dependent Schrodinger equation
$i\partial_{t}|\mathbf{v}(t)\rangle=-i\frac{\gamma}{2}|\mathbf{v}(t)\rangle+ie^{-\gamma t/2}\left(c_{+}\Gamma e^{\Gamma t}|u_{+}\rangle-c_{-}\Gamma e^{-\Gamma t}|u_{-}\rangle\right)=i\mathbf{L}_{\text{eff}}|\mathbf{v}(t)\rangle=i\mathbf{L}_{\text{eff}}e^{-\gamma t/2}\left(c_{+}e^{\Gamma t}|u_{+}\rangle+c_{-}e^{-\Gamma t}|u_{-}\rangle\right).$ (106)
with
$\mathbf{L}_{\text{eff}}|\mathbf{v}(t)\rangle$ $=e^{-\gamma t/2}\left(c_{+}e^{\Gamma t}\mathbf{L}_{\text{eff}}|u_{+}\rangle+c_{-}e^{-\Gamma t}\mathbf{L}_{\text{eff}}|u_{-}\rangle\right)$ (107)
$=e^{-\gamma t/2}\left[c_{+}e^{\Gamma t}\left(\Gamma-\frac{\gamma}{2}\right)|u_{+}\rangle+c_{-}e^{-\Gamma t}\left(-\Gamma-\frac{\gamma}{2}\right)|u_{-}\rangle\right]$
$=-\frac{\gamma}{2}e^{-\gamma t/2}(c_{+}e^{\Gamma t}|u_{+}\rangle+c_{-}e^{-\Gamma t}|u_{-}\rangle)+e^{-\gamma t/2}(c_{+}\Gamma e^{\Gamma t}|u_{+}\rangle-c_{-}\Gamma e^{-\Gamma t}|u_{-}\rangle)$
For $s<0$ we further have,
@@
$\langle u_{+}|\dot{\rho}(t)|u_{-}\rangle$ $=-\langle u_{+}|(\mathbf{L}_{\text{eff}}^{\dagger}\rho+\rho\mathbf{L}_{\text{eff}})|u_{-}\rangle=-[(\frac{\sqrt{-s}}{2}-\gamma/2)+(-\frac{\sqrt{-s}}{2}-\gamma/2)]\langle u_{+}|\rho(t)|u_{-}\rangle=\gamma\langle u_{+}|\rho(t)|u_{-}\rangle$
$\implies\langle u_{+}|\rho(t)|u_{-}\rangle=\langle u_{+}|\rho(0)|u_{-}\rangle e^{\gamma t},$
$\langle u_{-}|\dot{\rho}(t)|u_{+}\rangle$ $=-\langle u_{-}|(\mathbf{L}_{\text{eff}}^{\dagger}\rho+\rho\mathbf{L}_{\text{eff}})|u_{+}\rangle=-[(-\frac{\sqrt{-s}}{2}-\gamma/2)+(\frac{\sqrt{-s}}{2}-\gamma/2)]\langle u_{-}|\rho(t)|u_{+}\rangle=\gamma\langle u_{-}|\rho(t)|u_{+}\rangle$
$\implies\langle u_{-}|\rho(t)|u_{+}\rangle=\langle u_{-}|\rho(0)|u_{+}\rangle e^{\gamma t},$
@@ IX Appendix: Green’s Functions and Liouvillian Spectrum
The linear response of the system to a weak probe field is encoded in the local retarded Green’s function[22]. In the frequency domain, this can be expressed using the resolvent of the effective Liouvillian superoperator